How many grams of sodium benzoate (NaC7H5O2) is needed to make a 175.0mL buffer solution containing 0.15 M benzoic acid and a pH of 3.91?
1
Expert's answer
2020-04-18T07:06:28-0400

Solution.

"pH = pKa - lg(\frac{C(acid)}{C(base)})"

"Acid \ is \ C7H6O2"

"Base \ is \ C7H5O2^-"

"C(base) = C(acid) \times 10^{pH-pKa}"

C(base) = 0.075 M

"m(C7H5O2Na) = C(base) \times M(C7H5O2Na) \times V(solution)"

m(C7H5O2Na) = 1.90 g

Answer:

m(C7H5O2Na) = 1.90 g